Ali is 12 years younger than Mana. The sum of their ages is 80. How old is Mana?
Answer by tutoringisfun(17)

Good evening student! I hope this is helpful to you today.
First we need to label each variable
A=age of Ani M= age of Mana
from the problem we can deduce that A+M=80 and that A=M-12. We need to solve for either A or M. Since, A=M-12 was already given to me in the problem, I am going to use that for A. Next I substitute A=M-12 into the equation A+M=80
This gives the following:
M-12+M=80
to solve for M combine like terms and subtract 12 from each side, because we want M by itself.
2M=80+12
which can be rewritten as 2M=92
now to get X by itself we divide both sides by 2
2M/2=92/2 the two's on the left cancel out which gives us the following:
M=92/2
M=46
to calculate Ani's age we subtract 12 from 46 like so:
46-12=34. Therefore Ani is 34. We can check the answer by adding the two ages together. 46+34=80
I hope this has been helpful to you. If you desire long term assistance or need more help please contact me mbarre1@wgu.edu or visit my facebook page https://www.facebook.com/tutoringinquiry, or message me on instaedu.com